# On the maximal correlation of some stochastic processes

### v2

@@ 5 Analog and generalization of DKS inequality
$\sum_{(s,t):s\cap t\supset u}P(S=s,T=t)\sqrt{E\varphi_{s,u}^{2}(X_{u})}\sqrt{E\psi_{t,u}^{2}(X_{u})}\\ \leq R(S,T)\sqrt{\sum_{s:s\supset u}P(S=s)E\varphi_{s,u}^{2}(X_{u})}\sqrt{\sum_{t:t\supset u}P(T=t)E\psi_{t,u}^{2}(X_{u})}.$ (41)
By (5.2), (37), (5.2), (39), (41) and Cauchy-Schwarz inequality, we have that
@@
×$\quad\times\sum_{u\subset[n]}\sqrt{\sum_{s:s\supset u}P(S=s)E\varphi_{s,u}^{2}(X_{u})}\sqrt{\sum_{t:t\supset u}P(T=t)E\psi_{t,u}^{2}(X_{u})}$
@@
×$\quad\times\sqrt{\sum_{u\subset[n]}\sum_{s:s\supset u}P(S=s)E\varphi_{s,u}^{2}(X_{u})}\sqrt{\sum_{u\subset[n]}\sum_{t:t\supset u}P(T=t)E\psi_{t,u}^{2}(X_{u})}$
$=\max(R(S,T),\max(r_{u}:u\neq\emptyset))\sqrt{\operatorname{Var}\varphi(S,X_{S})}\sqrt{\operatorname{Var}\psi(T,X_{T})}.$ (42)
As $\varphi$ and $\psi$ is arbitrary, we conclude that
@@
By definition of $r_{u}$, for $u\subset\tilde{u}$, we have that $r_{u}\geq r_{\tilde{u}}$. Hence, we have that
@@
It remains to prove the converse inequality
@@
for some measurable function $h$ such that $Eh(X_{j})=0$, $\operatorname{Var}h(X_{j})<\infty$ and for $\alpha_{s}$ and $\beta_{t}$ achieving the best constant in the definition of $r_{\{j\}}$ (see (40)). It is strait-forward to check that for this particular choices of functions, we have that
$\operatorname{Cov}(\varphi(S,X_{S}),\psi(T,X_{T}))=r_{\{j\}}\sqrt{\operatorname{Var}\varphi(S,X_{S})\operatorname{Var}\psi(T,X_{T})}.$
@@
Hence, the converse inequality is proved.
When $S$ and $T$ are independent, we have $R(S,T)=0$. By writing $\widetilde{\alpha}_{s}=\sqrt{P(S=s)}\alpha_{s}$ and $\widetilde{\beta}_{t}=\sqrt{P(T=t)}\beta_{t}$, we see that $r_{\{j\}}$ is exactly the spectral radius of the rank-one matrix $(M_{st})_{s,t:j\in s\cap t}$, where $M_{st}=\sqrt{P(S=s)P(T=t)}$. Hence, we have that
@@
We prove Theorem 1.7 in this subsection. We will use the alternative definition (2) of the maximal correlation coefficient, the mathematical induction and the following lemma.
@@
Let $T$ be a uniform subset of $[n]=\{1,2,\ldots,n\}$ of size $n-1$. Given $T$, the random set $S$ is a uniform subset of $T$ of size $k$. Then, we have that
@@
Let $\varphi$ be a measurable function such that $E\varphi(T)=0$ and $\operatorname{Var}\varphi(T)<\infty$. Then, we have that
$E(\varphi(T)|S=s)=\sum_{i=1}^{n}\varphi(\{i\}^{c})P(T=\{i\}^{c}|S=s)$
@@
$E[(E(\varphi(T)|S)^{2}]$ $=\sum_{s\subset[n]}P(S=s)\left(\sum_{i=1}^{n}\varphi(\{i\}^{c})P(T=\{i\}^{c}|S=s)\right)^{2}$
@@
×$\quad\quad\quad\quad\quad\quad\times P(T=\{i\}^{c}|S=s)P(T=\{j\}^{c}|S=s)$
@@
=$=\left\{\begin{array}[]{ll}\frac{1}{(n-k)n},&i=j,\\ \frac{(n-k-1)}{n(n-1)(n-k)},&i\neq j.\end{array}\right.$
The only element outside of $T$ is uniformly distributed in $[n]$. Hence, the condition $E\varphi(T)=0$ and $\operatorname{Var}\varphi(T)<\infty$ is equivalent to
@@
and the value $\varphi(\{i\}^{c})$ is finite for each $i\in[n]$. Hence, by (2), we have that
$R^{2}(S,T)=\sup\{\sum_{i,j=1}^{n}A_{ij}\varphi(\{i\}^{c})\varphi(\{j\}^{c}):\sum_{i=1}^{n}\varphi(\{i\}^{c})=0,\sum_{i=1}^{n}(\varphi(\{i\}^{c}))^{2}=n\}.$
Note that $A=(A_{ij})_{i,j\in[n]}$ is a real symmetric matrix with positive elements. Hence, the eigenvalues of $A$ are real. By Perron-Frobenius theorem, $A$ has a simple eigenvalue $1/n$ and the moduli of other eigenvalues are strictly smaller than $1/n$. The eigenvector associated with $1/n$ is the column vector $\bm{1}=(1,1,\ldots,1)^{T}$. And the eigenvectors associated with other eigenvalues are perpendicular to $\bm{1}$. Hence, we get that
@@
where $I$ is the identity matrix. Hence, using a bit linear algebra, we find that $\lambda_{2}(A)=\frac{k}{n(n-1)(n-k)}$. Hence, $R(S,T)=\sqrt{n\lambda_{2}(A)}=\sqrt{\frac{k}{(n-1)(n-k)}}$. ∎
Next, we give the proof of Theorem 1.7 by induction on $n$. For $n=1$, $S$ and $T$ are independent. Hence, $R(S,T)=0$ and Theorem 1.7 holds. Suppose that Theorem 1.7 holds for $n\leq N$ and consider $n=N+1$. If $m=n$ or $m=0$, then $T$ is independent of $S$. Hence, $R(S,T)=0$ and Theorem 1.7 holds. Therefore, we may assume that $0<m<n$ in the following. Let $U$ be a random subset of $[n]$ with cardinality $n-1$ containing $T$. Moreover, assume that $U$ is uniform given $T$. In this way, we have $S\subset T\subset U$ and $(S,T,U)$ is uniform. Let $\varphi$ be a measurable function such that $E\varphi(S)=0$ and $E[(\varphi(S))^{2}]=1$. Since $(S,T,U)$ is Markov, i.e. $S$ and $U$ are conditionally independent given $T$, we have
$E(\varphi(S)|T)=E(\varphi(S)|T,U).$
@@
$g(U)=E(\varphi(S)|U).$
Then, we have that $E(f(S,U)|U)=0$ and $E(\varphi(S)|T,U)=g(U)+E(f(S,U)|T,U)$. Note that
$(E(\varphi(S)|T))^{2}$ $=(E(\varphi(S)|T,U))^{2}$
$=(g(U))^{2}+(E(f(S,U)|T,U))^{2}+2g(U)E(f(S,U)|T,U).$
Since $E(E(f(S,U)|T,U)|U)=E(f(S,U)|U)=0$, we have that
$E[(E(\varphi(S)|T))^{2}|U]=(g(U))^{2}+E[(E(f(S,U)|T,U))^{2}|U].$ (43)
Since $E(f(S,U)|U)=0$ and $\operatorname{Var}(f(S,U)|U)<\infty$, by (2), we have that
$E[(E(f(S,U)|T,U))^{2}|U]\leq(R(S,T|U))^{2}E[(f(S,U))^{2}|U],$ (44)
@@
$E[(E(\varphi(S)|T))^{2}|U]\leq\frac{k(n-1-m)}{m(n-1-k)}E[(\varphi(S))^{2}|U]+\left(1-\frac{k(n-1-m)}{m(n-1-k)}\right)(g(U))^{2}.$
@@
$E[(E(\varphi(S)|T))^{2}]\leq\frac{k(n-1-m)}{m(n-1-k)}E[(\varphi(S))^{2}]+\left(1-\frac{k(n-1-m)}{m(n-1-k)}\right)E[(g(U))^{2}].$ (47)
@@
Finally, combining (47) and (48), using the identity $\frac{k(n-1-m)}{m(n-1-k)}+(1-\frac{k(n-1-m)}{m(n-1-k)})\frac{k}{(n-1)(n-k)}=\frac{k(n-m)}{m(n-k)}$, we get that
$E[(E(\varphi(S)|T))^{2}]\leq\frac{k(n-m)}{m(n-k)}E[(\varphi(S))^{2}].$
Since this holds for all measurable $\varphi$ with $E\varphi(S)=0$ and $\operatorname{Var}\varphi(S)=1$, we conclude that
@@
To prove the converse inequality $R(S,T)\geq\sqrt{\frac{k(n-m)}{m(n-k)}}$, we take $\varphi(S)=1_{1\in S}$ and $\psi(T)=1_{1\in T}$. Then, we have that
@@
We will prove Corollary 1.8 in this subsection. In the proof, we use Theorem 1.6, Theorem 1.7 and Lemma 2.1 (submultiplicative property).
Firstly, since $\partial$ is different from the possible values of $X_{i}$, we have that
Let $U=S\cap T$. Then, we observe that $(S,X_{S}),(U,X_{U}),(T,X_{T})$ is Markov. Hence, by Lemma 2.1, we have that
$R((S,X_{S}),(T,X_{T}))\leq R((U,X_{U}),(S,X_{S}))R((U,X_{U}),(T,X_{T})).$ (49)
We explicitly calculate $R((U,X_{U}),(S,X_{S}))$ and $R((U,X_{U}),(T,X_{T}))$ in the following lemma:
@@
Recall that $(U,S)$ is uniformly distributed with the constraints that $|U|=\ell$, $|S|=k$ and $U\subset S$. We have that
$R((U,X_{U}),(S,X_{S}))=\sqrt{\frac{\ell}{k}}.$ (50)
Similarly, we have that
$R((U,X_{U}),(T,X_{T}))=\sqrt{\frac{\ell}{m}}.$ (51)
@@
We only give the proof of (50) as the proof of (51) is exactly the same. By Theorem 1.6, we have that
@@
where $r_{j}$ is the best constant $r$ for the following inequality
@@
$R(U,S)=\sqrt{\frac{\ell(n-k)}{k(n-\ell)}}\leq\sqrt{\frac{\ell}{k}}.$
@@
$r_{j}=\sqrt{\frac{\ell}{k}}.$ (53)
By performing the change of variables $\widetilde{\alpha}_{u}=\sqrt{P(U=u)}\alpha_{u}$ and $\widetilde{\beta}_{s}=\sqrt{P(S=s)}\beta_{s}$, we find that $r_{j}$ is exactly the spectral radius of the matrix $A=(A_{us})_{u,s:j\in u\cap s}$, where
@@
with the convention that $/0=0$ 0. Let $B=AA^{*}$, where $A^{*}$ is the transpose of $A$. Then, $B$ is real, symmetric and positively definite. Moreover, $r_{j}^{2}$ is exactly the maximal eigenvalue of $B$. Note that
@@
Then, we see that $B$ is similar to $C$, where
@@
with the convention that the conditional probability is zero if it is not well-defined. Hence, $r_{j}^{2}$ is equal to the maximal eigenvalue of $C$. Note that $P(j\in U|S=s)=\ell/k$ for $j\in s$ and $P(j\in S|U=u)=1$ for $j\in u$. Hence, we have that
@@
$=\frac{\ell}{k}\sum_{s:j\in s}P(S=s|U=u)$
$=\frac{\ell}{k}P(j\in S|U=u)$
$=\frac{\ell}{k}.$
Note that $C$ is a matrix with non-negative elements such that the sum of each row is the constant $\ell/k$. By Perron-Frobenius theorem, the maximal eigenvalue of $C$ is precisely $\ell/k$. Hence, we obtain that $r_{j}=\sqrt{\ell/k}$ and the proof is complete. ∎
By (49), (50) and (51), we obtain the upper bound
$R((S,X_{S}),(T,X_{T}))\leq\ell/\sqrt{km}.$
To prove the lower bound, we take a measurable function $h(x)$ such that $Eh(X_{j})=0$ and $\operatorname{Var}h(X_{j})=1$. Take $\varphi(S,X_{S})=1_{j\in S}h(X_{j})$ and $\psi(T,X_{T})=1_{j\in T}h(X_{j})$. Then, we get that
$R((S,X_{S}),(T,X_{T}))\geq\rho(\varphi(S,X_{S}),\psi(T,X_{T}))=\ell/\sqrt{km}.$
@@
Under the assumption of Theorem 1.6, by similar arguments as in Lemma 5.3, we have the following observation on $r_{j}$: If $P(j\in S|U=u)$ does not depend on $u$ when $j\in u$ and $P(j\in U|S=s)$ does not depend on $s$ when $j\in s$, then we have that
@@
We prove Theorem 1.9 in this subsection.
@@
by [MB07, Lemma 1] for the convolution identities for scores. Define
@@
where $u_{s}=\sum_{i\in s}x_{i}$. Then, we have that
$E[\varphi(S,X_{S})|T=t,X_{t}]=\left(\sum_{s\subset[n]}P(S=s|T=t)\lambda_{s}\right)\rho_{t}(U_{t})=\mu_{t}\rho_{t}(U_{t}).$
@@
$E[(E[\varphi(S,X_{S})|T,X_{T}])^{2}]=E\left[E\left[(E[\varphi(S,X_{S})|T,X_{T}])^{2}|T\right]\right]=\sum_{t\subset[n]}P(T=t)\mu_{t}^{2}I(U_{t}).$
@@
By (2), we have that
$E[(E[\varphi(S,X_{S})|T,X_{T}])^{2}]\leq R^{2}E[(\varphi(S,X_{S}))^{2}].$

### v3

@@ 5 Analogs and generalizations of DKS inequality
$\sum_{(s,t):s\cap t\supset u}P(S=s,T=t)\sqrt{E[\varphi_{s,u}^{2}(X_{u})]}\sqrt{E[\psi_{t,u}^{2}(X_{u})]}\\ \leq R(S,T)\sqrt{\sum_{s:s\supset u}P(S=s)E[\varphi_{s,u}^{2}(X_{u})]}\sqrt{\sum_{t:t\supset u}P(T=t)E[\psi_{t,u}^{2}(X_{u})]}.$ (42)
By (5.2), (38), (5.2), (40), (42) and the Cauchy-Schwarz inequality, we have that
@@
$\quad\times\sum_{u\subset[n]}\sqrt{\sum_{s:s\supset u}P(S=s)E[\varphi_{s,u}^{2}(X_{u})]}\sqrt{\sum_{t:t\supset u}P(T=t)E[\psi_{t,u}^{2}(X_{u})]}$
@@
$\quad\times\sqrt{\sum_{u\subset[n]}\sum_{s:s\supset u}P(S=s)E[\varphi_{s,u}^{2}(X_{u})]}\sqrt{\sum_{u\subset[n]}\sum_{t:t\supset u}P(T=t)E[\psi_{t,u}^{2}(X_{u})]}$
$=\max(R(S,T),\max(r_{u}:u\neq\emptyset))\sqrt{\operatorname{Var}(\varphi(S,X_{S}))}\sqrt{\operatorname{Var}(\psi(T,X_{T}))}.$ (43)
As $\varphi$ and $\psi$ are arbitrary, we conclude that
@@
By the definition of $r_{u}$, for $u\subset\tilde{u}$, we have that $r_{u}\geq r_{\tilde{u}}$. Hence, we have that
@@
It remains to prove the reverse inequality
@@
for some measurable function $h$ such that $E[h(X_{j})]=0$, $\operatorname{Var}(h(X_{j}))<\infty$ and for $\alpha_{s}$ and $\beta_{t}$ achieving the best constant in the definition of $r_{\{j\}}$ (see (41)). It is straightforward to check that for this particular choice of functions, we have that
$\operatorname{Cov}(\varphi(S,X_{S}),\psi(T,X_{T}))=r_{\{j\}}\sqrt{\operatorname{Var}(\varphi(S,X_{S}))\operatorname{Var}(\psi(T,X_{T}))}.$
@@
Hence, the reverse inequality is proved.
When $S$ and $T$ are independent, we have $R(S,T)=0$. By writing $\widetilde{\alpha}_{s}=\sqrt{P(S=s)}\alpha_{s}$ and $\widetilde{\beta}_{t}=\sqrt{P(T=t)}\beta_{t}$, we see that $r_{\{j\}}$ is precisely the spectral norm of the rank-one matrix $(M_{st})_{s,t:j\in s\cap t}$, where $M_{st}=\sqrt{P(S=s)P(T=t)}$. Hence, we have that
@@
In this subsection, we prove Theorem 1.6. We will use the alternative definition (2) of the maximal correlation coefficient, the mathematical induction and the following lemma.
@@
Let $T$ be a uniform subset of $n]=\{1,2,\ldots,n\}$ of size n1n-1. Given $T$, the random set $S$ is a uniform subset of $T$ of size $k$. Then we have that
@@
Let $\varphi$ be a measurable function such that $E[\varphi(T)]=0$ and $\operatorname{Var}(\varphi(T))<\infty$. Then we have that
$E[\varphi(T)|S=s]=\sum_{i=1}^{n}\varphi(\{i\}^{c})P(T=\{i\}^{c}|S=s)$
@@
$E[(E[\varphi(T)|S])^{2}]$ $=\sum_{s\subset[n]}P(S=s)\left(\sum_{i=1}^{n}\varphi(\{i\}^{c})P(T=\{i\}^{c}|S=s)\right)^{2}$
@@
$\quad\quad\quad\quad\quad\quad\times P(T=\{i\}^{c}|S=s)P(T=\{j\}^{c}|S=s)$
@@
$=\left\{\begin{array}[]{ll}\frac{1}{(n-k)n},&i=j,\\ \frac{(n-k-1)}{n(n-1)(n-k)},&i\neq j.\end{array}\right$.
The only element outside of $T$ is uniformly distributed in $[n]$. Hence, the condition $E[\varphi(T)]=0$ and $\operatorname{Var}(\varphi(T))<\infty$ is equivalent to
@@
and $\varphi(\{i\}^{c})\in\mathbb{R}$ for each $i\in[n]$. Hence, by (2), we have that
$R^{2}(S,T)=\sup\left\{\sum_{i,j=1}^{n}A_{ij}\varphi(\{i\}^{c})\varphi(\{j\}^{c}):\sum_{i=1}^{n}\varphi(\{i\}^{c})=0,\sum_{i=1}^{n}(\varphi(\{i\}^{c}))^{2}=n\right\}.$
Note that $A=(A_{ij})_{i,j\in[n]}$ is a real symmetric matrix with positive elements. Hence, the eigenvalues of $A$ are real. By the Perron-Frobenius theorem, $A$ has a simple eigenvalue $1/n$, and the moduli of other eigenvalues are strictly less than $1/n$. The eigenvector associated with $1/n$ is the column vector $\bm{1}=(1,1,\ldots,1)^{T}$. And the eigenvectors associated with other eigenvalues are perpendicular to $\bm{1}$. Hence, we get that
@@
where $I$ is the identity matrix. Hence, using a bit of linear algebra, we find that $\lambda_{2}(A)=\frac{k}{n(n-1)(n-k)}$. Hence, $R(S,T)=\sqrt{n\lambda_{2}(A)}=\sqrt{\frac{k}{(n-1)(n-k)}}$. ∎
Next, we give the proof of Theorem 1.6 by induction on $n$. For $n=1$, $S$ and $T$ are independent. Hence, $R(S,T)=0$ and Theorem 1.6 holds. Suppose that Theorem 1.6 holds for $n\leq N$, and consider $n=N+1$. If $m=n$ or $m=0$, then $T$ is independent of $S$. Hence, $R(S,T)=0$ and Theorem 1.6 holds. Therefore, we may assume that $0<m<n$ in the following. Let $U$ be a random subset of $[n]$ with cardinality n1n-1 that contains $T$. Moreover, assume that $U$ is uniform given $T$. In this way, we have $S\subset T\subset U$, and $(S,T,U)$ is uniformly distributed. Let $\varphi$ be a measurable function such that $E[\varphi(S)]=0$ and $E[(\varphi(S))^{2}]=1$. Since $(S,T,U)$ is Markov, we have that
$E[\varphi(S)|T]=E[\varphi(S)|T,U].$
@@
$g(U)=E[\varphi(S)|U].$
Then we have that $E[f(S,U)|U]=0$ and $E[\varphi(S)|T,U]=g(U)+E[f(S,U)|T,U]$. Note that
$(E[\varphi(S)|T])^{2}$ $=(E[\varphi(S)|T,U])^{2}$
$=(g(U))^{2}+(E[f(S,U)|T,U])^{2}+2g(U)E[f(S,U)|T,U].$
Since $E[E[f(S,U)|T,U]|U]=E[f(S,U)|U]=0$, we have that
$E[(E[\varphi(S)|T])^{2}|U]=(g(U))^{2}+E[(E[f(S,U)|T,U])^{2}|U].$ (44)
Since $E[f(S,U)|U]=0$ and $\operatorname{Var}(f(S,U)|U)<\infty$, by (2), we have that
$E[(E[f(S,U)|T,U])^{2}|U]\leq(R(S,T|U))^{2}E[(f(S,U))^{2}|U],$ (45)
@@
$E[(E[\varphi(S)|T])^{2}|U]\leq\frac{k(n-1-m)}{m(n-1-k)}E[(\varphi(S))^{2}|U]+\left(1-\frac{k(n-1-m)}{m(n-1-k)}\right)(g(U))^{2}.$
@@
$E[(E[\varphi(S)|T])^{2}]\leq\frac{k(n-1-m)}{m(n-1-k)}E[(\varphi(S))^{2}]+\left(1-\frac{k(n-1-m)}{m(n-1-k)}\right)E[(g(U))^{2}].$ (48)
@@
Finally, combining (48) and (49), using the identity
$\frac{k(n-1-m)}{m(n-1-k)}+\left(1-\frac{k(n-1-m)}{m(n-1-k)}\right)\frac{k}{(n-1)(n-k)}=\frac{k(n-m)}{m(n-k)},$
we get that
$E[(E[\varphi(S)|T])^{2}]\leq\frac{k(n-m)}{m(n-k)}E[(\varphi(S))^{2}].$
Since this holds for all measurable $\varphi$ with $E[\varphi(S)]=0$ and $\operatorname{Var}(\varphi(S))=1$, we conclude that
@@
To prove the reverse inequality $R(S,T)\geq\sqrt{\frac{k(n-m)}{m(n-k)}}$, we take $\varphi(S)=1_{1\in S}$ and $\psi(T)=1_{1\in T}$. Then we have that
@@
In this subsection, we prove Corollary 1.7. In the proof, we use Theorem 1.5, Theorem 1.6 and Lemma 2.1 (submultiplicative property).
Let $(S,T)$ be uniformly distributed with the constraints that
$S\subset[n],T\subset[n],|S|=m,|T|=n-\ell\text{ and }|S\cap T|=m-\ell.$
Suppose that $(S,T)$ is independent of $X_{1},X_{2},\ldots,X_{n}$. Let $U=S\cap T$. Since $X_{1},X_{2},\ldots,X_{n}$ are i.i.d., $(\sum_{i\in S}\delta_{X_{i}},\sum_{i\in T}\delta_{X_{i}})$ has the same distribution as $(\sum_{i=1}^{m}\delta_{X_{i}},\sum_{i=\ell+1}^{n}\delta_{X_{i}})$. Moreover, $\sum_{i\in S}\delta_{X_{i}}$ and $\sum_{i\in T}\delta_{X_{i}}$ are conditionally independent given $\sum_{i\in U}\delta_{X_{i}}$. Hence, by Lemma 2.1, we have that
$R\left(\sum_{i=1}^{m}\delta_{X_{i}},\sum_{i=\ell+1}^{n}\delta_{X_{i}}\right)$ $=R\left(\sum_{i\in S}\delta_{X_{i}},\sum_{i\in T}\delta_{X_{i}}\right)$
$\leq R\left(\sum_{i\in S}\delta_{X_{i}},\sum_{i\in U}\delta_{X_{i}}\right)R\left(\sum_{i\in U}\delta_{X_{i}},\sum_{i\in T}\delta_{X_{i}}\right)$
$\leq R((U,X_{U}),(S,X_{S}))R((U,X_{U}),(T,X_{T})).$ (50)
We explicitly calculate $R((U,X_{U}),(S,X_{S}))$ in the following lemma.
@@
Consider independent non-degenerate random variables $X_{1},X_{2},\ldots,X_{n}$ taking values in a general measurable space. Let $(U,S)$ be uniformly distributed with the constraints that $|U|=a$, $|S|=b$ and $U\subset S\subset[n]$. Suppose that $(U,S)$ is independent of $X_{1},X_{2},\ldots,X_{n}$. Then we have that
$R((U,X_{U}),(S,X_{S}))=\sqrt{a/b}.$ (51)
Remark 5.2.
Lemma 5.3 is an extension of the Dembo-Kagan-Shepp inequality for independent random variables $X_{1},X_{2},\ldots,X_{n}$ with possibly different distributions.
@@
By Theorem 1.5, we have that
@@
where $r_{j}$ is the best constant $r$ in the following inequality
@@
$R(U,S)=\sqrt{\frac{a(n-b)}{b(n-a)}}\leq\sqrt{\frac{a}{b}}.$
@@
$r_{j}=\sqrt{\frac{a}{b}}.$ (53)
By changing the variables $\widetilde{\alpha}_{u}=\sqrt{P(U=u)}\alpha_{u}$ and $\widetilde{\beta}_{s}=\sqrt{P(S=s)}\beta_{s}$, we find that $r_{j}$ is precisely the spectral norm of the matrix $A=(A_{us})_{u,s:j\in u\cap s}$, where
@@
with the convention that $0/0=0$. Let $B=AA^{*}$, where $A^{*}$ denotes the transpose of $A$. Then $B$ is real, symmetric and positively definite. Moreover, $r_{j}^{2}$ is exactly the maximal eigenvalue of $B$. Note that
@@
Then we see that $B$ is similar to $C$, where
@@
with the convention that the conditional probability is zero if it is not well-defined. Hence, $r_{j}^{2}$ is equal to the maximal eigenvalue of $C$. Note that $P(j\in U|S=s)=a/b$ for $j\in s$ and $P(j\in S|U=u)=1$ for $j\in u$. Hence, we have that
@@
$=\frac{a}{b}\sum_{s:j\in s}P(S=s|U=u)$
$=\frac{a}{b}P(j\in S|U=u)$
$=\frac{a}{b}.$
Note that $C$ is a matrix with non-negative elements such that the sum of each row is the constant $a/b$. By the Perron-Frobenius theorem, the maximal eigenvalue of $C$ is precisely $a/b$. Thus, $r_{j}=\sqrt{a/b}$, and the proof is complete. ∎
By (5.4) and (51), we obtain the upper bound
$R\left(\sum_{i=1}^{m}\delta_{X_{i}},\sum_{i=\ell+1}^{n}\delta_{X_{i}}\right)\leq\sqrt{\frac{m-\ell}{m}}\sqrt{\frac{m-\ell}{n-\ell}}=\frac{m-\ell}{\sqrt{m(n-\ell)}}.$
For the lower bound, by the Dembo-Kagan-Shepp-Yu equality ([Yu08, Theorem 4.1]), we have that
$R\left(\sum_{i=1}^{m}\delta_{X_{i}},\sum_{i=\ell+1}^{n}\delta_{X_{i}}\right)\geq R\left(\sum_{i=1}^{m}\varphi(X_{i}),\sum_{i=\ell+1}^{n}\varphi(X_{i})\right)=\frac{m-\ell}{\sqrt{m(n-\ell)}}$
for some non-degenerate real-valued measurable function $\varphi$.
@@
Under the assumption of Theorem 1.5, using similar arguments as in Lemma 5.3, we have the following observation about $r_{j}$: If $P(j\in S|U=u)$ does not depend on $u$ when $j\in u$ and $P(j\in U|S=s)$ does not depend on $s$ when $j\in s$, then we have that
@@
In this subsection, we prove Theorem 1.8 .
@@
by [MB07, Lemma 1] (i.e. the convolution identity for scores). Define
@@
where $u_{s}=\sum_{i\in s}x_{i}$. Then we have that
$E[\varphi(S,X_{S})|T=t,U_{t}]=\left(\sum_{s\subset[n]}P(S=s|T=t)\lambda_{s}\right)\rho_{t}(U_{t})=\mu_{t}\rho_{t}(U_{t}).$
@@
$E[(E[\varphi(S,X_{S})|T,U_{T}])^{2}]=E\left[E\left[(E[\varphi(S,X_{S})|T,U_{T}])^{2}|T\right]\right]=\sum_{t\subset[n]}P(T=t)\mu_{t}^{2}I(U_{t}).$
@@
By (2) and the fact that $(T,U_{T})$ is a measurable function of $(T,X_{T})$, we have that
$E[(E[\varphi(S,X_{S})|T,U_{T}])^{2}]\leq R^{2}((T,U_{T}),(S,X_{S}))E[(\varphi(S,X_{S}))^{2}]\leq R^{2}E[(\varphi(S,X_{S}))^{2}].$